OpenStudy (anonymous):

Two circles have areas of 16(pi) and 25(pi). Find the ratio of their circumferences.

OpenStudy (eyust707):

First off you should memorize and know this: \[\pi = {c \over d}\]

OpenStudy (eyust707):

c is circumference d is diameter

OpenStudy (eyust707):

Also memorize and know: \[Area \space of \space a \space circle = \pi r^2\]

OpenStudy (eyust707):

r is the radius

OpenStudy (anonymous):

Sorry There was mistake in the last step..

OpenStudy (eyust707):

So we know the areas... with that information we can find the radius of both circles using the formula I posted

OpenStudy (eyust707):

Lets plug the first one in

OpenStudy (eyust707):



OpenStudy (anonymous):

4/5 as ratio of there ara is 16/25=R^2/R^2 thus ratios of perimeter would be r/R then taking aquare root we have r/R=4/5

OpenStudy (eyust707):

okay sweet the pi's cancel

OpenStudy (eyust707):